# Effects of Flagellar Morphology on Swimming Performance and Directional Control in Microswimmers

### v1

@@ Abstract
In a fluid environment, flagellated microswimmers propel themselves by rotating their flagella. The morphology of these flagella significantly influences forward speed, swimming efficiency, and directional stability, which are critical for their survival. This study begins by simulating the three-dimensional motion trajectories of microswimmers to analyze their kinematic characteristics. The simulation results demonstrate that microswimmers can actively adjust their forward direction by modifying the orientation of their flagella. We subsequently perform numerical simulations to visualize the flow fields generated by a microswimmer and examine the hydrodynamic interaction between the cell body and the flagella, focusing on their impacts on forward speed and swimming efficiency. Our findings indicate that forward speed and swimming efficiency are closely related to the filament radius, pitch angle, and contour length of the flagella, while the yaw angle of locomotion is determined by the helix radius and contour length of the flagella. Our findings indicate that the pitch angle for maximum forward speed is slightly smaller than that for maximum swimming efficiency, suggest that microswimmers can effectively alternate between states of maximum forward speed and maximum swimming efficiency by fine-tuning their pitch angle and adapting to varying ecological conditions. These morphological characteristics of microswimmers may result from species competition and natural selection. This research establishes an optimized model for microswimmers, providing valuable insights for the design of enhanced microrobots tailored to specific applications.
@@ I Introduction
Flagellated bacteria are a crucial group of microorganisms playing a vital role in natural ecosystems primarily due to their unique swimming mechanisms and exceptional motility. These bacteria are typically composed of a cell body, flagella, hooks, and motors, with the beating or rotation of the flagella serving as the primary mechanism for self-propulsion Silverman and Simon (1974); Lauga and Powers (2009); Lauga (2016). Research on bacterial motility has advanced not only the fields of microbiology and physics, but also provided significant insights into ecological functions Guasto, Rusconi, and Stocker (2012). The morphological characteristics of these bacteria have evolved over time to adapt to diverse living environments. This evolution directly influences their survival strategies and enhances their motility through their specific morphological structures Roszak and Colwell (1987); Young (2006); Mitchell and Kogure (2006); Justice et al. (2008); Spagnolie and Lauga (2011); van Teeseling, de Pedro, and Cava (2017).
@@
Marine bacteria exhibit rapid forward speeds, high swimming efficiencies, and strong directional responses to their ecological environment. These traits enable them to quickly react to chemical gradients, which ensures efficient nutrient acquisition and reproduction Maki et al. (2000); Stocker and Seymour (2012). Previous studies have shown that the morphological characteristics of flagella, such as filament radius, pitch angle, and contour length, significantly influence bacterial forward speed, swimming efficiency, and motility stability Spagnolie and Lauga (2011); Purcell (1997); Li and Tang (2006); Chattopadhyay et al. (2006); Najafi et al. (2018); Nguyen and Graham (2018). Collectively, these factors determine the competitiveness of bacteria in complex environments, thereby influencing their survival and reproduction Fauci and Dillon (2006).
@@
In recent years, rapid advancements in experimental techniques and numerical simulation capabilities have enabled researchers to explore the physical mechanisms of bacterial motility in greater depth Silverman and Simon (1974); Bianchi, Saglimbeni, and Di Leonardo (2017); Thawani and Tirumkudulu (2018); Zöttl and Yeomans (2019); Colin et al. (2021). The locomotion of bacteria in fluids generates specific flow fields that not only influence the swimming trajectories of microswimmers but also significantly affect the movement of surrounding microorganisms. Understanding these flow characteristics is essential for a comprehensive understanding of the motion mechanisms of microorganisms and the hydrodynamic interactions among their components Zöttl and Yeomans (2019); Drescher et al. (2010, 2011); Mathijssen, Pushkin, and Yeomans (2015). Variations in parameters such as flagellar filament radius, helix radius, pitch angle, and axial length can significantly impact the forward speed, swimming efficiency, precession angle (also referred to as the wobbling angle), and yaw angle of bacteria Shum, Gaffney, and Smith (2010); Son, Menolascina, and Stocker (2016); Mousavi, Gompper, and Winkler (2020); Tang et al. (2020); Tokárová et al. (2021). Therefore, establishing an appropriate model of single-flagellated microswimmers and conducting systematic simulation studies is crucial for revealing their optimal morphological characteristics.
@@
In-depth research on the motion mechanisms of microswimmers provides valuable insights into their biological characteristics while serving as a key source of inspiration for the design of new microrobots. With the rapid advancements in microbiology and engineering technologies, researchers have conducted more comprehensive investigations into the motion mechanisms and optimal morphologies of single-flagellated microswimmers Nelson, Kaliakatsos, and Abbott (2010); Temel and Yesilyurt (2015); Ng et al. (2021). Microrobots, as an emerging technology, demonstrate wide-ranging application potential across various fields, including biomedicine, environmental monitoring, and micro-manipulation Palagi and Fischer (2018); Zhou et al. (2021). By harnessing the motion mechanisms of single-flagellated bacteria, researchers can design more efficient and flexible microrobots that adapt to complex working environments Huang et al. (2019).
@@
In this study, we systematically investigate the dynamic and kinematic characteristics of a helical flagellated microswimmer model in a viscous fluid using numerical simulation methods. By employing the twin multipole moment (TMM) method to calculate the system’s resistance matrix, we effectively account for the hydrodynamic interaction between the cell body and the flagella Jeffrey and Onishi (1984); Liu, Chen, and Zhang (2025). We adjust several morphological parameters of the microswimmer model, including the filament radius, helix radius, pitch angle, and contour length of the flagella, to assess its motion trajectory, flow field, forward speed, swimming efficiency, and yaw angle. These simulations help us determine the optimal parameters for the microswimmer design. Collectively, these factors influence the competitiveness of bacteria in complex environments, which subsequently impacts their survival and reproductive success in various contexts.
@@
The structure of this paper is organized as follows: In Sec. II, we establish the microswimmer simulation model and analyze its kinematics and swimming efficiency. In Sec. III. A, we simulate the trajectory of the microswimmer and analyze its kinematic characteristics. Following that, in Sec. III. B, we examine the flow field generated by the microswimmer during swimming, specifically analyzing how the hydrodynamic interaction between the cell body and the flagella affects its forward speed and swimming efficiency. In Sec. III. C, we conduct a detailed numerical simulation to assess the impact of varying filament radius, pitch angle, and contour length on the forward speed and swimming efficiency of the microswimmer, ultimately identifying the optimal filament radius, pitch angle, and contour length. Additionally, we analyze how variations in contour length and helix radius influence forward speed and swimming efficiency, integrating the yaw angle and helical diameter into our examination of locomotion trajectory at different helical radii to identify the optimal helix radius. Finally, in Sec. IV, we summarize the main findings of our investigation.
@@ II.1 Model and Methods
where $\mathbf{U}^{\infty}$ and $\bm{\Omega}^{\infty}$ denote the ambient flow fields, while $\mathbf{U}-\mathbf{U}^{\infty}$ and $\bm{\Omega}-\bm{\Omega}^{\infty}$ represent the relative velocities of the $N$ spheres in a $3N$-dimensional space relative to the ambient flow. The resistance matrix of the system, denoted as $\mathcal{R}$, can be calculated using the twin multipole moment (TMM) Liu, Chen, and Zhang (2025). The reciprocal relation that relates the velocities of the microswimmer and the external forces and torques in the ambient flow fields are related by the mobility matrix $\mathcal{M}$, which can be expressed as
@@
Flagellated microswimmers generally consist of a cell body, flagella, motors, and hooks Silverman and Simon (1974); Lauga and Powers (2009); Lauga (2016); Nord et al. (2022). In most numerical simulations of rigid microswimmers, the effects of hooks are typically neglected Abbott et al. (2009); Acemoglu and Yesilyurt (2014); Vizsnyiczai et al. (2020). In this study, we assume that the specific morphology of the hooks can be neglected and choose to omit them. The microswimmer is modeled as a rigid helical flagellum attached to a spherical cell body with a radius $R_{b}$, as illustrated in Fig. 1. The helical flagellum is characterized by its helix radius $R$, pitch $\lambda$, contour length $\Lambda$, and filament radius $a$. Its axial length is given by $L=\Lambda\cos\theta$, where the pitch angle $\theta$ satisfies $\tan\theta=2\pi R/\lambda$. The center of the cell body is denoted by $\mathbf{r}_{b}$, while the coordinate of the connection point between the flagellum and the cell body is represented as $\mathbf{r}_{0}$. The associated material frame at the connection point $\mathbf{r}_{0}$ is defined as $\{\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{D}_{3}\}$, where $\mathbf{D}_{1}$ is the unit vector along the flagellar axis pointing towards the cell body. We define the unit vector indicating the rotation direction of the microswimmer’s motor as $\mathbf{e}_{m}=\frac{\mathbf{r}_{0}-\mathbf{r}_{b}}{|\mathbf{r}_{0}-\mathbf{r}_{b}|}$. In the laboratory coordinate system, the centerline positions of the left-handed helical flagellum are given by:
@@
$\begin{split}&\mathbf{F}_{b}+\sum_{i=1}^{N}\mathbf{F}_{t}^{i}=0,\\ &\mathbf{T}_{b}+\sum_{i=1}^{N}\mathbf{T}_{t}^{i}+\sum_{i=1}^{N}\left(\mathbf{r}^{i}-\mathbf{r}_{b}\right)\times\mathbf{F}_{t}^{i}=0.\end{split}$ (6)
where $\mathbf{F}_{t}$ and $\mathbf{T}_{t}$ represent, respectively, the forces and torques exerted by the spheres on the fluid along the flagellum, and $N$ denotes the total number of spheres.
@@ II.2 Kinematic Analysis
Analyzing the kinematics of microswimmers is crucial for assessing their performance. Since both the cell body and the flagellum of the microswimmer are rigid, the rotation of the flagellum induces non-axisymmetric behavior, resulting in precessional motion within the fluid and leading to a helical trajectory. The forward direction of the microswimmer is defined by the rotation of the flagellum, expressed as $\mathbf{e}_{t}=\bm{\Omega}_{t}/|\bm{\Omega}_{t}|$ Constantino et al. (2016); Benhal et al. (2021); Kamdar et al. (2022). The forward speed $U_{f}$ is given by $U_{f}=-\mathbf{U}_{b}\cdot\mathbf{e}_{t}$, while the speed perpendicular to the forward direction is $U_{v}=\sqrt{U_{b}^{2}-U_{f}^{2}}$, where $U_{b}$ is the translational velocity of the cell body. Consequently, the pitch of the helical trajectory is given by $H=\frac{2\pi}{|\bm{\Omega}_{t}|}U_{f}$, and the helical diameter is given by $D=\frac{2}{|\bm{\Omega}_{t}|}U_{v}$.
@@ II.3 Swimming Efficiency
Swimming efficiency is an important parameter characterizing the performance of a microswimmer. The propulsive force acting in the forward direction is expressed as $F_{f}=-\mathbf{F}_{b}\cdot\mathbf{e}_{t}$. The propulsive efficiency $\epsilon$ is defined as the ratio of the propulsive power output to the rotary power input provided by the motor, which refers to the power expended to rotate the flagellum Purcell (1997); Chattopadhyay et al. (2006); Spagnolie and Lauga (2011); Acemoglu and Yesilyurt (2014):
@@
where $\mathbf{T}_{b}$ represents the torque exerted on the cell body by the motor, and $\bm{\Omega}_{m}$ denotes the rotational velocity of the motor.
@@ III.1 Trajectories of Microswimmers
Figure 2 illustrates the trajectories of microswimmers with varying contour lengths, pitch angles, and helix radii over a duration of $0.03$ $\mathrm{s}$. It is clear that as the contour length increases, the forward speed of the microswimmer also increases. Furthermore, an increase in the helix radius $R$ corresponds to a larger helical diameter $D$ of the trajectories. Notably, when the pitch angle is $\pi/5$, the forward speed exceeds that of the pitch angles $\pi/6$ and $\pi/4$. The microswimmer’s trajectory forms a cylindrical helix, with the forward direction $\mathbf{e}_{t}$ defined as a fixed unit vector. From Fig. 2, it is apparent that the morphological parameters of different microswimmers lead to variations in forward speeds and helical radii of the trajectories, indicating the presence of an optimal flagellar morphology that enables the microswimmer to achieve maximum forward speed.
@@
Figure 3: (a) Schematic diagram of the microswimmer in its initial state. (b), (c), and (d) illustrate the cones formed by the rotation in different directions around the forward axis over time, corresponding to various angles $\psi$ between the motor’s rotation direction and the flagellar axis.
@@
When a microswimmer moves through a fluid, several specific directional vectors must be carefully considered. Fig. 3 illustrates the time-dependent changes of these unit vectors during the locomotion of the microswimmers. In Fig. 3(a), the initial model of the microswimmer is depicted, with $\mathbf{e}_{m}$ representing the direction of the motor’s rotation and $\mathbf{e}_{a}$ indicating the direction of the flagellar axis, separated by an angle $\psi$. Fig. 3(b) shows the variations of these directions over time when $\psi=0^{\circ}$. In this case, the forward direction remains unchanged, while the purple arrow denotes the spin direction of the cell body, which rotates around the forward axis to form a green cone. The angle between the spin direction and the forward direction is referred to as the precession angle, while the angle between the flagellar axis and the forward direction is minimized, indicating that these two directions are nearly aligned. Fig. 3(c) illustrates the directional variations when $\psi=10^{\circ}$. In this scenario, the green cone represents the spin direction of the cell body rotating around the forward direction, the dark blue cone indicates the motor’s rotation direction, and the gold cone denotes the flagellar axis rotating around the forward direction. Similarly, Fig. 3(d) presents the variations of these directions when $\psi=20^{\circ}$.
@@
Comparing Figs. 3(b)-(d), it is evident that the precession angle is the largest, followed by the angle between the motor’s rotation direction and the forward direction. Conversely, the angle between the flagellar axis direction and the forward direction is the smallest. When the motor’s rotation direction aligns with the flagellar axis, the angle between the forward direction and the flagellar axis approaches a minimum, which we define as the yaw angle. Since the direction of the flagellum is controllable, the forward direction of the microswimmer can be adjusted by manipulating the orientation of the flagellum.
@@
Figure 4: Closed trajectory of a microswimmer’s four-step locomotion involves alternating both the flagellar direction and the initial phase at each step, while adjusting the motor’s rotation direction every two steps.
@@
From Fig. 3, it is clear that the direction of the flagellum is crucial for controlling the locomotion direction of the microswimmer. The angle between the flagellar axis and the forward direction, referred to as the yaw angle plays a critical role in determining the navigation capabilities of the flagellum. To intuitively investigate the swimming trajectory of the microswimmer in the fluid, we simulate its locomotion in four steps, as illustrated in Fig. 4. In all four steps, the angle is set at $\psi=30^{\circ}$. In the first step, the motor rotates along the negative $x$-axis. In the second step, the flagellar axis is symmetrically reflected about the $x$-axis, and the initial phase of the flagellum is increased by half a cycle $\pi$ to maintain the morphological symmetry of the microswimmer, while the motor’s rotation direction remains unchanged. During the third step, the flagellar axis is symmetrically reflected about the $y$-axis, and the initial phase is again increased by half a cycle $\pi$, with the motor’s rotation direction now aligned along the positive $x$-axis. In the fourth step, the flagellar axis is symmetrically reflected about the $x$-axis. Once again, the initial phase is increased by half a cycle $\pi$, while the motor’s rotation direction continues along the positive $x$-axis. Each step lasts for $3$ $\mathrm{s}$, resulting in a total duration of 12 $\mathrm{s}$, after which the microswimmer returns to its initial position. The enlarged portion indicates that the locomotion trajectory of the microswimmer follows a helical path. This simulation demonstrates that the forward direction of the microswimmer is governed by the flagellum, with changes in the forward direction corresponding to alterations in the flagellar direction Lauga (2016); Kühn et al. (2017). Therefore, navigation can be achieved by adjusting the flagellar direction, and the entire locomotion process follows a helical trajectory (Supplementary Movie).

### v2

@@ Abstract
In a fluid environment, flagellated microswimmers propel themselves by rotating their flagella. The morphology of these flagella significantly influences forward speed, swimming efficiency, and directional stability, which are critical for their survival. This study begins by simulating the three-dimensional motion trajectories of microswimmers to analyze their kinematic characteristics. The simulation results demonstrate that microswimmers can actively adjust their forward direction by modifying the orientation of their flagella. We subsequently perform numerical simulations to visualize the flow fields generated by a microswimmer and examine the hydrodynamic interactions between the cell body and the flagella, focusing on their impacts on forward speed and swimming efficiency. We conclude that forward speed and swimming efficiency are closely related to the filament radius, pitch angle, and contour length of the flagella, while the yaw angle of locomotion is determined by the helix radius and contour length of the flagella. We conclude that the pitch angle for maximum forward speed is slightly smaller than that for maximum swimming efficiency, which suggests that microswimmers can effectively alternate between states of maximum forward speed and maximum swimming efficiency by fine-tuning their pitch angle and adapting to varying ecological conditions. These morphological characteristics of microswimmers may result from species competition and natural selection. This research establishes an optimized model for microswimmers, providing valuable insights for the design of enhanced microrobots tailored to specific applications.
@@ I Introduction
Flagellated bacteria are a crucial group of microorganisms playing a vital role in natural ecosystems primarily due to their unique swimming mechanisms and exceptional motility. These bacteria are typically composed of a cell body, flagella, hooks, and motors, with the beating or rotation of the flagella serving as the primary mechanism for self-propulsion Silverman and Simon (1974); Lauga and Powers (2009); Lauga (2016). Research on bacterial motility has not only advanced the fields of microbiology and physics, but also provided significant insights into ecological functions Guasto, Rusconi, and Stocker (2012). The morphological characteristics of these bacteria have evolved over time to adapt to diverse living environments. This evolution directly influences their survival strategies and enhances their motility through their specific morphological structures Roszak and Colwell (1987); Young (2006); Mitchell and Kogure (2006); Justice et al. (2008); Spagnolie and Lauga (2011); van Teeseling, de Pedro, and Cava (2017).
@@
Marine bacteria exhibit rapid forward speeds, high swimming efficiencies, and strong directional responses to their ecological environment. These traits enable them to quickly react to chemical gradients, ensuring efficient nutrient acquisition and reproduction Maki et al. (2000); Stocker and Seymour (2012). Previous studies have shown that the morphological characteristics of flagella, such as filament radius, pitch angle, and contour length, significantly influence bacterial forward speed, swimming efficiency, and motility stability Spagnolie and Lauga (2011); Purcell (1997); Li and Tang (2006); Chattopadhyay et al. (2006); Najafi et al. (2018); Nguyen and Graham (2018). Collectively, these factors determine the competitiveness of bacteria in complex environments and thereby influence their survival and reproduction Fauci and Dillon (2006).
@@
In recent years, rapid advancements in experimental techniques and numerical simulation capabilities have enabled researchers to explore the physical mechanisms of bacterial motility in greater depth Silverman and Simon (1974); Bianchi, Saglimbeni, and Di Leonardo (2017); Thawani and Tirumkudulu (2018); Zöttl and Yeomans (2019); Colin et al. (2021). The locomotion of bacteria in fluids generates specific flow fields that not only influence the swimming trajectories of microswimmers but also significantly affect the movement of surrounding microorganisms. Understanding these flow characteristics is essential for a comprehensive understanding of the motion mechanisms of microorganisms and the hydrodynamic interactions among their components Zöttl and Yeomans (2019); Drescher et al. (2010, 2011); Mathijssen, Pushkin, and Yeomans (2015). Variations in parameters such as flagellar filament radius, helix radius, pitch angle, and axial length can significantly impact the forward speed, swimming efficiency, precession angle (also referred to as the wobbling angle), and yaw angle of bacteria Shum, Gaffney, and Smith (2010); Son, Menolascina, and Stocker (2016); Mousavi, Gompper, and Winkler (2020); Tang et al. (2020); Tokárová et al. (2021). Therefore, establishing an appropriate model of single-flagellated microswimmers and conducting systematic simulation studies are crucial for revealing their optimal morphological characteristics.
@@
In-depth research on the motion mechanisms of microswimmers provides valuable insights into their biological characteristics and serves as a key source of inspiration for the design of new microrobots. With the rapid advancements in microbiology and engineering technologies, researchers have conducted more comprehensive investigations into the motion mechanisms and optimal morphologies of single-flagellated microswimmers Nelson, Kaliakatsos, and Abbott (2010); Temel and Yesilyurt (2015); Ng et al. (2021). Microrobots, as an emerging technology, demonstrate wide-ranging application potential across various fields, including biomedicine, environmental monitoring, and micro-manipulation Palagi and Fischer (2018); Zhou et al. (2021). By harnessing the motion mechanisms of single-flagellated bacteria, researchers can design more efficient and flexible microrobots that can adapt to complex working environments Huang et al. (2019).
@@
In this study, we systematically investigate the dynamic and kinematic characteristics of a helical flagellated microswimmer model in a viscous fluid using numerical simulation methods. By employing the Twin Multipole Moment (TMM) method to calculate the resistance matrix of the system, we effectively account for the hydrodynamic interactions between the cell body and the flagella. The TMM method considers both near-field and far-field interactions, making it suitable for high-resolution studies of the hydrodynamics of microswimmers Jeffrey and Onishi (1984); Liu, Chen, and Zhang (2025). We adjust several morphological parameters of the microswimmer model, including the filament radius, helix radius, pitch angle, and contour length of the flagella, to assess its motion trajectory, flow field, forward speed, swimming efficiency, and yaw angle. These simulations help us determine the optimal parameters for the microswimmer design. Collectively, these factors influence the competitiveness of bacteria in complex environments, which subsequently impacts their survival and reproductive success in various contexts.
@@
The structure of this paper is organized as follows: In Sec. II, we establish the microswimmer simulation model and analyze its kinematics and swimming efficiency. In Sec. III. A, we simulate the trajectory of the microswimmer and analyze its kinematic characteristics. Following that, in Sec. III. B, we examine the flow field generated by the microswimmer during swimming, specifically analyzing how the hydrodynamic interactions between the cell body and the flagella affects its forward speed and swimming efficiency. In Sec. III. C, we conduct a detailed numerical simulation to assess the impact of varying filament radius, pitch angle, and contour length on the forward speed and swimming efficiency of the microswimmer, ultimately identifying the optimal filament radius, pitch angle, and contour length. Additionally, we analyze how variations in contour length and helix radius influence forward speed and swimming efficiency, integrating the yaw angle and helical diameter into our examination of locomotion trajectories at different helical radii to identify the optimal helix radius. Finally, in Sec. IV, we summarize the main findings of our investigation.
@@ II.1 Model and Methods
where $\mathbf{U}^{\infty}$ and $\bm{\Omega}^{\infty}$ denote the ambient flow fields, while $\mathbf{U}-\mathbf{U}^{\infty}$ and $\bm{\Omega}-\bm{\Omega}^{\infty}$ represent the relative velocities of the $N$ spheres in a $3N$-dimensional space relative to the ambient flow. The resistance matrix of the system, denoted as $\mathcal{R}$, can be calculated using TMM Liu, Chen, and Zhang (2025). The reciprocal relation that relates the velocities of the microswimmer to the external forces and torques in the ambient flow fields is expressed using the mobility matrix $\mathcal{M}$, which can be expressed as
@@
Flagellated microswimmers generally consist of a cell body, flagella, motors, and hooks Silverman and Simon (1974); Lauga and Powers (2009); Lauga (2016); Nord et al. (2022). In most numerical simulations of rigid microswimmers, the effects of hooks are typically neglected Abbott et al. (2009); Acemoglu and Yesilyurt (2014); Vizsnyiczai et al. (2020). In this study, we assume that the specific morphology of the hooks can be neglected and choose to omit them. The microswimmer is modeled as a rigid helical flagellum attached to a spherical cell body with a radius $R_{b}$, as illustrated in Fig. 1. The helical flagellum is characterized by its helix radius $R$, pitch $\lambda$, contour length $\Lambda$, and filament radius $a$. Its axial length is given by $L=\Lambda\cos\theta$, where the pitch angle $\theta$ satisfies $\tan\theta=2\pi R/\lambda$. The center of the cell body is denoted by $\mathbf{r}_{b}$, while the coordinate of the connection point between the flagellar axis and the cell body is represented as $\mathbf{r}_{0}$. The associated material frame at the connection point $\mathbf{r}_{0}$ is defined as $\{\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{D}_{3}\}$, where $\mathbf{D}_{1}$ is the unit vector along the flagellar axis pointing towards the cell body. We define the unit vector indicating the rotation direction of the microswimmer’s motor as $\mathbf{e}_{m}=\frac{\mathbf{r}_{0}-\mathbf{r}_{b}}{|\mathbf{r}_{0}-\mathbf{r}_{b}|}$. In the laboratory coordinate system, the centerline positions of the left-handed helical flagellum are given by:
@@
$\begin{split}&\mathbf{F}_{b}+\sum_{i=1}^{N-1}\mathbf{F}_{t}^{i}=0,\\ &\mathbf{T}_{b}+\sum_{i=1}^{N-1}\mathbf{T}_{t}^{i}+\sum_{i=1}^{N-1}\left(\mathbf{r}^{i}-\mathbf{r}_{b}\right)\times\mathbf{F}_{t}^{i}=0.\end{split}$ (6)
where $\mathbf{F}_{b}$ and $\mathbf{T}_{b}$ represent the force and torque exerted by the cell body on the fluid, respectively. Similarly, $\mathbf{F}_{t}$ and $\mathbf{T}_{t}$ denote the forces and torques exerted by the flagellum on the fluid. The number of spheres on the flagellum is $N-1$, while the total number of spheres in the microswimmer model is $N$.
@@ II.2 Kinematic Analysis
Analyzing the kinematics of microswimmers is crucial for assessing their performance. The rigid structure of both the cell body and the flagellum causes non-axisymmetric behavior during rotation, resulting in precessional motion and a helical trajectory. The forward direction of the microswimmer is defined by the rotation of the flagellum, expressed as $\mathbf{e}_{t}=\bm{\Omega}_{t}/|\bm{\Omega}_{t}|$ Constantino et al. (2016); Benhal et al. (2021); Kamdar et al. (2022). The forward speed $U_{f}$ is given by $U_{f}=-\mathbf{U}_{b}\cdot\mathbf{e}_{t}$, while the speed perpendicular to the forward direction is $U_{v}=\sqrt{U_{b}^{2}-U_{f}^{2}}$, where $\mathbf{U}_{b}$ is the translational velocity of the cell body. Consequently, the pitch of the helical trajectory is given by $H=\frac{2\pi}{|\bm{\Omega}_{t}|}U_{f}$, and the helical diameter is given by $D=\frac{2}{|\bm{\Omega}_{t}|}U_{v}$.
@@ II.3 Swimming Efficiency
Swimming efficiency is an important parameter that characterizes the performance of a microswimmer. The propulsive force acting in the forward direction is expressed as $F_{f}=-\mathbf{F}_{b}\cdot\mathbf{e}_{t}$. The propulsive efficiency $\epsilon$ is defined as the ratio of the propulsive power output to the rotary power input provided by the motor, which refers to the power expended to rotate the flagellum Purcell (1997); Chattopadhyay et al. (2006); Spagnolie and Lauga (2011); Acemoglu and Yesilyurt (2014):
@@
where $\mathbf{T}_{b}$ represents the torque exerted on the cell body by the motor, and $\bm{\Omega}_{m}$ denotes the rotational velocity of the motor. The various parameters of the model for the microswimmer are summarized in Table. 1.
Table 1: The various parameters in the numerical model and their values.
Notation Description Value
$\mu$ Dynamic viscosity $1.0$ $\mu\mathrm{g}\mathrm{/}\mathrm{(}\mu\mathrm{m}\cdot\mathrm{s}\mathrm{)}$
$R_{b}$ Radius of cell body $1.0$ $\mu\mathrm{m}$
$f$ Rotation frequency of motor $100$ $\mathrm{H}\mathrm{z}$
$\mathbf{e}_{a}$ Flagellar axis direction $\mathbf{e}_{a}=\mathbf{D}_{1}$
$\mathbf{e}_{m}$ Rotation direction of motor $\mathbf{e}_{m}=\frac{\mathbf{r}_{0}-\mathbf{r}_{b}}{|\mathbf{r}_{0}-\mathbf{r}_{b}|}$
$\bm{\Omega}_{m}$ Angular velocity of motor $2\pi f\mathbf{e}_{m}$
$\bm{\Omega}_{b}$ Angular velocity of cell body Variable
$\bm{\Omega}_{t}$ Angular velocity of flagellum $\bm{\Omega}_{t}=\bm{\Omega}_{b}+\bm{\Omega}_{m}$
$-\mathbf{e}_{t}$ Forward direction $-\bm{\Omega}_{t}/|\bm{\Omega}_{t}|$
$\psi$ Angle between flagellar axis and rotation direction of motor $\arccos(-\mathbf{e}_{m}\cdot\mathbf{e}_{a})$
$\beta$ Yaw angle $\arccos(-\mathbf{e}_{t}\cdot\mathbf{e}_{a})$
$U_{f}$ Forward speed $-\mathbf{U}_{b}\cdot\mathbf{e}_{t}$
$U_{v}$ Speed perpendicular to forward direction $\sqrt{U_{b}^{2}-U_{f}^{2}}$
$H$ Pitch of helical trajectory $\frac{2\pi}{|\bm{\Omega}_{t}|}U_{f}$
$D$ Helical diameter of trajectory $\frac{2}{|\bm{\Omega}_{t}|}U_{v}$
@@ III.1 Trajectories of Microswimmers
Figure 2 illustrates the trajectories of microswimmers with varying contour lengths, pitch angles, and helix radii over a duration of $0.03$ $\mathrm{s}$. It is clear that as the contour length increases, the forward speed of the microswimmer also increases. Furthermore, an increase in the helix radius $R$ corresponds to a larger helical diameter $D$ of the trajectories. Notably, when the pitch angle is $\pi/5$, the forward speed exceeds that of the pitch angles $\pi/6$ and $\pi/4$. The microswimmer’s trajectory forms a cylindrical helix, with the forward direction $-\mathbf{e}_{t}$ is a fixed unit vector. From Fig. 2, it is apparent that the morphological parameters of different microswimmers lead to variations in forward speeds and helical radii of the trajectories, indicating the presence of an optimal flagellar morphology that enables the microswimmer to achieve maximum forward speed.
@@
Figure 3: (a) Schematic diagram of the microswimmer in its initial state. (b), (c), and (d) illustrate the cones formed by rotation in different directions around the forward axis over time, corresponding to various angles $\psi$ between the motor’s rotation direction and the flagellar axis.
@@
When a microswimmer moves through a fluid, several specific directional vectors must be carefully considered. Fig. 3 illustrates the time-dependent changes in these unit vectors during the locomotion of the microswimmers. In Fig. 3(a), the initial model of the microswimmer is depicted, with $\mathbf{e}_{m}$ representing the direction of the motor’s rotation and $\mathbf{e}_{a}$ indicating the direction of the flagellar axis, separated by an angle $\psi$. Fig. 3(b) shows the variations of these directions over time when $\psi=0^{\circ}$. In this case, the forward direction remains unchanged, while the purple arrow denotes the spin direction of the cell body, which rotates around the forward axis to form a green cone. The angle between the spin direction and the forward direction is referred to as the precession angle, while the angle between the flagellar axis and the forward direction approaches zero, indicating that these two directions are nearly aligned. Fig. 3(c) illustrates the directional variations when $\psi=10^{\circ}$. In this scenario, the green cone represents the spin direction of the cell body rotating around the forward direction, the dark blue cone indicates the motor’s rotation direction, and the gold cone denotes the flagellar axis rotating around the forward direction. Similarly, Fig. 3(d) presents the variations of these directions when $\psi=20^{\circ}$.
@@
As shown in Figs. 3(b)-(d), the precession angle is the largest, followed by the angle between the motor’s rotation and the forward direction. Conversely, the angle between the flagellar axis direction and the forward direction is the smallest. When the motor’s rotation direction aligns with the flagellar axis, the angle between the forward direction and the flagellar axis approaches a minimum, which we define as the yaw angle. Since the direction of the flagellum is controllable, the forward direction of the microswimmer can be adjusted by manipulating the orientation of the flagellum.
@@
Figure 4: Closed trajectory of a microswimmer’s four-step locomotion involves alternating both the flagellar direction and the initial phase at each step, while adjusting the motor’s rotation direction every two steps. Multimedia available online.
@@
From Fig. 3, it is clear that the direction of the flagellum is crucial for controlling the locomotion direction of the microswimmer. The angle between the flagellar axis and the forward direction, referred to as the yaw angle, plays a critical role in determining the navigation capabilities of the flagellum. To intuitively investigate the swimming trajectory of the microswimmer in the fluid, we simulate its locomotion in four steps, as illustrated in Fig. 4 (Multimedia available online). In all four steps, the angle is set at $\psi=30^{\circ}$. In the first step, the motor rotates along the negative $x$-axis. In the second step, the flagellar axis is symmetrically reflected about the $x$-axis, and the initial phase of the flagellum is increased by half a cycle ($\pi$) to maintain the morphological symmetry of the microswimmer, while the motor’s rotation direction remains unchanged. During the third step, the flagellar axis is symmetrically reflected about the $y$-axis, and the initial phase is again increased by half a cycle ($\pi$), with the motor’s rotation direction now aligned along the positive $x$-axis. In the fourth step, the flagellar axis is symmetrically reflected about the $x$-axis once again, and the initial phase is increased by half a cycle ($\pi$), while the motor’s rotation direction continues along the positive $x$-axis. Each step lasts for $3$ $\mathrm{s}$, resulting in a total duration of 12 $\mathrm{s}$, after which the microswimmer returns to its initial position. The enlarged portion indicates that the locomotion trajectory of the microswimmer follows a helical path. This simulation demonstrates that the forward direction of the microswimmer is governed by the flagellum, with changes in the forward direction corresponding to alterations in the flagellar direction Lauga (2016); Kühn et al. (2017). Therefore, navigation can be achieved by adjusting the flagellar direction, and the entire locomotion process follows a helical trajectory.